# BondBERT: What we learn when assigning sentiment in the bond market

### v1

@@ Abstract
Bond markets respond differently to macroeconomic news compared to equity markets, yet most sentiment models, including FinBERT, are trained primarily on general financial or equity news data. This mismatch is important because bond prices often move in the opposite direction to economic optimism, making general or equity-based sentiment tools potentially misleading. In this paper, we introduce BondBERT, a transformer-based language model fine-tuned on bond-specific news. BondBERT can act as the perception and reasoning component of a financial decision-support agent, providing sentiment signals that integrate with forecasting models. It is a generalisable framework for adapting transformers to low-volatility, domain-inverse sentiment tasks by compiling and cleaning 30,000 UK bond market articles (2018–2025) for training, validation, and testing. We compare BondBERT’s sentiment predictions against FinBERT, FinGPT, and Instruct-FinGPT using event-based correlation, up/down accuracy analyses, and LSTM forecasting across ten UK sovereign bonds. We find that BondBERT consistently produces positive correlations with bond returns, achieves higher alignment and forecasting accuracy than the three baseline models, with lower normalised RMSE and higher information coefficient. These results demonstrate that domain-specific sentiment adaptation better captures fixed income dynamics, bridging a gap between NLP advances and bond market analytics.
@@ 1 INTRODUCTION
Bond markets are fundamental to global finance, yet they remain comparatively underexplored in artificial intelligence research. Sentiment analysis applies techniques to quantify the tone or emotional score of textual data, converting qualitative news, reports, or social-media posts into numerical scores. By providing a quantitative measure of market mood, sentiment analysis has become an important tool for modelling and predicting trends across financial markets (Souza et al.,, 2015; Luna et al.,, 2015; Pagolu et al.,, 2016; Shabani et al.,, 2020; Li et al.,, 2020). This approach has been shown to be particularly insightful in fixed-income analysis, where a growing body of research demonstrates a clear link between sentiment and the behaviour of bond markets (Piñeiro-Chousa et al.,, 2021; Bartov et al.,, 2023; Erlwein-Sayer,, 2018; Bakar and Hamzah,, 2022).
However, there are no tools specifically introduced for the fixed income market. This gap is critical because bond markets react differently to macroeconomic news, where positive economic developments often lead to higher yields and lower bond prices, a dynamic which is opposite to equities. As a result, sentiment models trained on general or equity focused corpora can misinterpret bond-specific signals.
Despite evidence linking sentiment to the bond market, no transformer model has been fine-tuned specifically for bond market sentiment. In this paper we introduce BondBERT, a transformer-based sentiment model fine-tuned on 17,000 bond-related news articles spanning from 2018–2022. We aim to show how domain-adaptive fine-tuning can extend transformers into domains with inverse sentiment-price relationships, small effect sizes, and noisy signals.
We benchmark BondBERT against FinBERT, FinGPT, and Instruct-FinGPT across three tasks: (i) correlation with bond log returns on sentiment shock days, (ii) up/down accuracy in event studies, (iii) graphical representation of sentiment and bond market returns, and (iv) bond price predictions using a LSTM model.
Thus, we try to address two main research questions: (i) Does a bond-specific transformer (BondBERT) produce sentiment more aligned with bond price movements than other baselines and (ii) Does BondBERT sentiment carry predictive power for future bond prices? Our results show that BondBERT performs better than the baselines in understanding the market and has a stronger predictive power, with a lower normalised RMSE and a higher information coefficient. [^1] Pearson correlation between predicted and actual prices.
@@
The remainder of this paper is organised as follows: Section 2 reviews previous research related to financial sentiment analysis, Section 3 describes our dataset and experimental setup, and Section 4 covers the architecture of BondBERT and training procedure. Finally, Sections 5 and 6 present the results of our research and outline future research directions, respectively.
@@ 2 Literature review
More recently, transformer-based models such as FinBERT have emerged as a benchmark for financial sentiment analysis. Introduced by Araci, (2019), FinBERT extends BERT through additional domain-adaptive pretraining on financial corpora including Reuters, Yahoo Finance, corporate reports, and earnings transcripts (Du et al.,, 2024). Comparative studies consistently show FinBERT outperforms both the base BERT model and traditional machine learning approaches, largely due to its contextual embeddings and financial domain adaptation (Dataloop,, 2025). Araci’s original work demonstrated significant accuracy gains even with limited fine-tuning data (Araci,, 2019), and subsequent studies confirm its robustness and better performance relative to other large language models across various financial sentiment datasets (Yang et al.,, 2020; Mert et al.,, 2025; Du et al.,, 2024; Arshadi et al.,, 2025).
@@
While FinBERT has been widely applied to equity market forecasting (Zhang,, 2025; Gu et al.,, 2024; Kulkarni et al.,, 2022; Singh et al.,, 2024; Sreehitha et al.,, 2024; Rajesh et al.,, 2024; Chen et al.,, 2025), its use in bond markets has been far more limited, with only a handful of studies exploring this application (Makara,, 2024; Liu et al.,, 2024; Talebi et al.,, 2025; Sontakke and Patil,, 2024). Importantly, no variant of FinBERT has yet been trained or adapted specifically for bond-related text. In this work, we address this gap by employing the FinBERT model introduced by Yang et al., (2020) model and further adapting it for the analysis of bond market sentiment, as we will explain in more detail in Section 4.
@@
The development of FinGPT represents one of the first systematic efforts to create an open-source financial large language model. The framework was introduced as a pipeline that automates large-scale financial data collection, with lightweight adaptation methods such as LoRA/QLoRA, and proposes reinforcement learning with stock prices for alignment (Liu et al.,, 2023). Its architecture was described in four layers (data, adaptation, evaluation, application), providing a foundation for customisable financial AI research (Yang et al.,, 2023). Building on the above, an instruction-tuning benchmark for FinGPT was proposed, showing that relatively small models (7 billion parameters) fine-tuned with finance-specific instructions can match or even outperform larger untuned models (Wang et al.,, 2023). This highlighted the need for instruction-tuned financial LLMs for sentiment analysis. Thus, we start seeing studies using FinGPT for predictive modelling. Liang et al., (2024) developed a dissemination-aware FinGPT for stock prediction, where company news is grouped to capture the dissemination breadth, while context is added to the prompts. They achieved an improvement of 8% in the prediction of short-term movement over LLM using raw news. Moreover, Zhou and Mehra, (2025) proposed an end-to-end trading framework that combines FinGPT sentiment pipelines, illustrating the feasibility of incorporating LLMs into trading workflows.
@@
Another popular tool, Instruct-FinGPT, explores instruction tuning as a method to adapt general-purpose LLMs in finance. This is achieved by converting small supervised financial sentiment datasets into instruction-style data and fine-tuning LLaMA-7B, as shown in Zhang et al., (2023). Their models outperformed FinBERT and zero-shot ChatGPT in sentiment classification, showing improved performance in context-sensitive inputs. Wang et al., (2023) further supported this, confirming the need of instruction tuned models in finance. Moreover, based on systematic evaluations of instruction-tuned LLMs across finance it was found that instruction tuning consistently improves performance, enabling smaller models to close the gap with larger ones (Fatemi et al.,, 2025). However, we do see limitations when it comes beyond sentiment analysis, such as long-form document parsing (e.g., annual reports and 10-Ks) or valuation modelling (Li et al.,, 2023; Jadhav and Mirza,, 2025).
@@
Our work connects to research on decision-support systems, since BondBERT can be viewed as a perceptual agent module embedded within larger multi-agent financial architectures, where different agents handle forecasting, risk assessment, and trading.
@@ 3 Experimental setup
We utilise two primary data sources, bond market transactional data and financial news articles. The transactional bond market data was provided by Propellant.digital, a Software as a Service (SaaS) company that has developed a platform that aggregates and standardises trading data across markets, serving as a working prototype for a consolidated tape in the bond market. [^2] https://propellant.digital/ The data was extracted from their Big Query client.api and the subset we used covers all trading days from January 2022 to January 2025. To ensure representative market movements, bonds were filtered for sufficient liquidity and volatility, and daily prices were transformed into logarithmic returns to allow for better analysis. Thus, the dataset includes price information for the 10 most liquid UK sovereign bonds in that period. It should be noted that UK corporate bonds were excluded from the analysis due to their illiquidity, which prevented us from conducting predictive experiments with the sentiment analysis tools.
Table 1 presents the average and median prices of the 10 UK instruments. Prices are denoted as a percentage of the notional (face) value, meaning, for example, that a price of 98.5 corresponds to 98.5% of the bond’s face value. The table highlights the relatively close price range among these instruments. This is further supported by the distribution of the prices in Figure 1, where we observe that each instrument shows low volatilities, with the majority of the log returns around 0.
Table 1: Instrument prices (2022–2025).
Instrument ID Average Price Median Price
GB0004893086 106.14 102.42
GB0032452392 105.39 100.66
GB00B24FF097 108.72 104.87
GB00B3Y1JG82 102.95 99.23
GB00B52WS153 108.01 103.52
GB00BDRHNP05 92.73 92.05
GB00BFX0ZL78 92.29 91.37
GB00BMGR2916 70.91 68.55
GB00BYZW3G56 95.29 94.63
GB00BZB26Y51 78.24 74.72
Furthermore, Figure 2 illustrates the monthly liquidity of the instruments, providing insight into their trading activity levels. Understanding liquidity is important for our later predictions as bonds with higher trading activity provide more reliable and representative price data, while low liquidity instruments can exhibit irregular price movements that can distort predictive performance and the evaluation of the two sentiment analysis tools.
Figure 2: Average number of trades for the 10 UK instruments from 2022 to 2025.
@@
All preprocessing scripts, labelling prompts, and model fine-tuning configurations will be released on GitHub for reproducibility.
@@ 4 Methodology
To address the absence of a bond-specific sentiment tool, we developed BondBERT by further adapting a financial domain language model for the fixed-income context. Specifically, we start from the FinBERT model introduced by Yang et al., (2020), which is domain pretrained on a large scale financial text (e.g., SEC filings and Reuters news) but not fine-tuned on labeled sentiment datasets. We fine-tuned this model on our bond-specific sentiment dataset to create BondBERT. For comparison, we use ProsusAI/FinBERT (ProsusAI,, 2020) as a baseline, which is widely used in prior work and is fine-tuned on the Financial PhraseBank dataset (Malo et al.,, 2014) (general financial news sentiment) but not specialised for bonds.
@@
The training corpus for BondBERT comprised approximately 17,000 bond-related news articles spanning 2018–2021, internationally, covering macroeconomic, fiscal, and policy domains relevant to bond markets. [^4] Topics included are “macroeconomic trends”, “macroeconomic outlook”, “bond market”, “global economic growth”, “economic outlook”, “economic uncertainty”, “interest rates”, “monetary policy”, “global risk sentiment”, “financial market volatility”, “European bond market”, “energy crisis”, “global inflation”, “global GDP slowdown”, “macroeconomic report”, “bond market”, “investor sentiment”, “sanctions and trade wars”, and “geopolitical tensions”. These articles were first cleaned for duplicates, irrelevant topics, and insufficient content length, then aligned to market dates. Sentiment labels were generated by employing GPT-4.1-nano via the OpenAI API and evaluating different prompts designed to capture the expected impact of each article on bond prices. [^5] We evaluated different GPT models, such as GPT-4.1 and GPT-4.1-mini, but due to similar results, we opted for GPT-4.1-nano, which is more computationally efficient.
@@
We used GPT-4.1-nano to generate sentiment labels efficiently, testing prompt formats until we identified one that produced consistent and financially relevant outputs. This was critical because general-purpose prompts often misinterpret fixed income market signals (e.g., interpreting “higher interest rates” as positive rather than negative for bond prices). However, we recognise that this approach introduces potential label noise and bias due to the model’s prior training data and prompt sensitivity. To mitigate this, we conducted prompt engineering experiments and cross-checked a subset of labels against known market reactions, though future work should incorporate human annotated or hybrid validation to improve model accuracy.
@@
We constructed a set of twelve representative financial news snippets, four clearly negative, four clearly positive, and four neutral, based on real bond market reporting. These examples included diverse contexts such as central bank rate hikes, inflation reports, bond issuance announcements, and fiscal policy updates. Each candidate prompt was evaluated on its alignment with expected bond price reactions and its consistency across similar examples. Example articles and their sentiment scores are given below:
Positive
“Bank of England announces unexpected rate cut to support growth”: 0.8
Neutral
“UK government releases annual debt issuance schedule in line with forecasts”: 0
Negative
“UK inflation surges to highest level in 30 years, raising rate hike fears”: -0.9
@@
The best performing prompt was concise and framed the model as a bond market analyst. This framing improved label consistency, ensuring that positive scores reflected bond price increases rather than general economic optimism. Finally, we asked the model to assign a sentiment score between 1-1 (strongly negative for bond prices) and +1+1 (strongly positive), with 0 indicating neutral. The final prompt, which was deliberately kept concise for better accuracy, was:
@@
Using this prompt, we processed 17,000 articles for training, almost 2,000 for validation, and 2,000 for testing, in order to generate continuous sentiment scores in the range [1,1][-1,1]. These scores served as the training labels for BondBERT’s supervised fine-tuning stage. These sentiment scores were merged with the article text using titles as unique keys to ensure accurate label alignment.
We initially considered further fine-tuning FinBERT on bond-specific sentiment data. However, because FinBERT’s prior training was optimised for general financial related sentiment, we expected additional fine-tuning might conflict with the existing scores and introduce noise rather than improve alignment with bond price movements. For this reason, we developed BondBERT as a separate model, pre-trained on financial text but fine-tuned exclusively on bond market data. Sentiment scores were categorised into three categories, negative, neutral, and positive, and stratified undersampling was applied to mitigate class imbalance by creating equal sample sizes across the categories.
@@
Table 2: BondBERT configuration parameters
@@
Both FinBERT and BondBERT output probabilities for three sentiment classes: negative, neutral, and positive. To enable comparison and subsequent correlation analysis with bond returns, we convert these probabilities into a continuous sentiment score $s\in[-1,1]$ using the Normalised Difference Index (NDI), as in Borovkova and Zhang, (2022). The NDI is defined as:
@@
where $s$ is the score of the sentiment, p(positive) is the probability of the text being positive and p(negative) the probability of the text being negative. This transformation is particularly useful in our context because it explicitly removes the neutral class and amplifies the polarity of the sentiment signal. By focusing only on the positive and negative components, the NDI score more directly captures information that can move the market. This improves alignment when correlating sentiment with bond price returns.
@@
We used Instruct-FinGPT in its 13B parameter version, accessed through the Hugging Face inference endpoint. [^7] https://huggingface.co/FinGPT/Instruct-FinGPT-13B Unlike the FinGPT setup, no additional LoRA adapters were attached, since the model is already instruction-tuned on financial sentiment. Most importantly, Instruct-FinGPT expects inputs formatted as task-specific instructions followed by a financial text passage. This is so it can produce more reliable sentiment outputs compared to zero-shot prompting. For a fair comparison, we used the same version we used for the BondBERT prompt, broken down in smaller, individual sentences, adding more guidelines to avoid chattiness:
@@
Finally, it is worth clarifying that because both FinGPT and Instruct-FinGPT are not stand-alone architectures, we had to use LLaMA-2 as their underlying backbone. These models were released as adaptations of the LLaMA-2 family (7B and 13B), with LoRA adapters provided for FinGPT and instruction-tuned checkpoints for Instruct-FinGPT specifically designed for this backbone. Although newer open-source backbones such as LLaMA-3 and LLaMA-4 are released, the FinGPT resources currently available are not compatible with them. Since our goal was to compare BondBERT against other financial models, we included FinGPT and Instruct-FinGPT in the evaluation despite their dependence on LLaMA-2.
One drawback, however, of using GPT-based annotation is the potential risk of peek-ahead bias. Since GPT models are trained on large post-event corpora, they may inadvertently leverage knowledge of subsequent market outcomes when labelling historical articles. In our current setup we did not impose strict temporal framing in the prompt, so we cannot fully rule out this risk. However, the prompt was designed to request the immediate expected impact of each article on bond prices, which partly reduces the chance of future knowledge being incorporated. Future work should address this more rigorously by enforcing time-consistent prompting or complementing synthetic labels with human annotation. This applies not only to GPT-4.1-nano, but also to Instruct-FinGPT and any LLM-based annotator.
(a) FinBERT
(b) BondBERT
(c) FinGPT
(d) Instruct-FinGPT
Figure 4: Directional accuracy of sentiment tools across the four models relative to bond returns.

### v2

@@ Abstract
Bond markets respond differently to macroeconomic news compared to equity markets, yet most sentiment models are trained primarily on general financial or equity news data. However, bond prices often move in the opposite direction to economic optimism, making general or equity-based sentiment tools potentially misleading. We introduce BondBERT, a transformer-based language model fine-tuned on bond-specific news. BondBERT can act as the perception and reasoning component of a financial decision-support agent, providing sentiment signals that integrate with forecasting models. We propose a generalisable framework for adapting transformers to low-volatility, domain-inverse sentiment tasks by compiling and cleaning 30,000 UK bond market articles (2018–2025). BondBERT’s sentiment predictions are compared against FinBERT, FinGPT, and Instruct-FinGPT using event-based correlation, up/down accuracy analyses, and LSTM forecasting across ten UK sovereign bonds. We find that BondBERT consistently produces positive correlations with bond returns, and achieves higher alignment and forecasting accuracy than the three baseline models. These results demonstrate that domain-specific sentiment adaptation better captures fixed income dynamics, bridging a gap between NLP advances and bond market analytics.
@@ 1 INTRODUCTION
Bond markets are fundamental to global finance, yet remain comparatively under-explored in artificial intelligence research. Sentiment analysis quantifies the tone or emotional score of textual data, converting qualitative news, reports, or social-media posts into numerical scores. It has long been known that media sentiment can predict market trading volumes (Tetlock,, 2007), and recent works demonstrate that sentiment analysis can predict bond market returns (Bartov et al.,, 2023; Isakin and Pu,, 2023; Piñeiro-Chousa et al.,, 2021).
Several transformer-based models, such as FinBERT (Araci,, 2019) and FinGPT (Wang et al.,, 2023), have been fine-tuned on financial datasets. Yet, while it is known that bond markets react differently to equity markets on news (Isakin and Pu,, 2023), there are no models fine-tuned specifically for bond market sentiment. To address this gap, we introduce BondBERT, a transformer-based sentiment model fine-tuned on 17,000 bond-related news articles covering 2018–2022. We aim to show how domain-adaptive fine-tuning can extend transformers into domains with inverse sentiment-price relationships, small effect sizes, and noisy signals.
We benchmark BondBERT against FinBERT, FinGPT, and Instruct-FinGPT across four tasks: (i) correlation with bond returns on sentiment shock days, (ii) event-study directional accuracy, (iii) graphical representation between sentiment and returns, and (iv) bond price prediction using an LSTM model.
Thus, we address two questions: (i) Does a bond-specific transformer (BondBERT) produce sentiment more aligned with bond price movements than other baselines? and (ii) Does BondBERT sentiment carry predictive power for future bond prices? Results show that BondBERT performs better than the benchmarks, with stronger predictive power, lower normalised RMSE, and higher information coefficient.
@@
The remainder of this paper is organised as follows: Section 2 reviews previous research related to financial sentiment analysis, Section 3 describes our dataset and experimental setup, and Section 4 covers the architecture of BondBERT and training procedure. Finally, Sections 5 and 6 present our results and outline future research directions, respectively.
@@ 2 Literature review
More recently, transformer-based models such as FinBERT have emerged as a benchmark for financial sentiment analysis. Introduced by Araci, (2019), FinBERT extends BERT through additional domain-adaptive pretraining on financial corpora including Reuters, Yahoo Finance, corporate reports, and earnings transcripts (Du et al.,, 2024). Comparative studies show that FinBERT outperforms both the base BERT model and traditional machine learning approaches, largely due to its contextual embeddings and financial domain adaptation. [^1] https://dataloop.ai/library/model/prosusai˙finbert/ Araci’s original work demonstrated significant accuracy gains even with limited fine-tuning data (Araci,, 2019), and subsequent studies confirm its robustness and better performance relative to other large language models across various financial sentiment datasets (Yang et al.,, 2020; Kang and Choi,, 2025; Du et al.,, 2024; Huang et al.,, 2023).
@@
While FinBERT has been widely applied to equity market forecasting (Fan et al.,, 2024; Gu et al.,, 2024; Cicekyurt and Bakal,, 2025; Zhang,, 2025), its use in bond markets has been limited (Makara,, 2024; Liu et al.,, 2025). Importantly, no variant of FinBERT has yet been trained or adapted specifically for bond-related text. In this work, we address this gap by employing the FinBERT model introduced by Yang et al., (2020), and adapt it to the analysis of sentiment of bond markets (see details in Section 4).
@@
The development of FinGPT represents one of the first systematic efforts to create an open-source financial large language model. The framework introduced an automated pipeline for large-scale financial data collection, lightweight adaptation methods such as LoRA/QLoRA, and reinforcement learning with stock prices for alignment (Liu et al.,, 2023). Its four-layer architecture provides a foundation for customisable financial AI research (Yang et al.,, 2023). Building on the above, an instruction-tuning benchmark for FinGPT was proposed, showing that relatively small models (7 billion parameters) fine-tuned with finance-specific instructions can match or even outperform larger untuned models (Wang et al.,, 2023). Thus, recent studies used FinGPT for predictive modelling. Liang et al., (2024) developed a dissemination-aware FinGPT for stock prediction, where company news is grouped to capture dissemination breadth, while context is added to the prompts, achieving an improvement of 8% in the prediction of short-term movement over LLM using raw news. Moreover, Zhou and Mehra, (2025) proposed an end-to-end trading framework that combines FinGPT sentiment pipelines, illustrating the feasibility of incorporating LLMs into trading workflows.
@@
Another popular tool, Instruct-FinGPT, applies instruction tuning to adapt general-purpose LLMs in finance by converting small supervised sentiment datasets into instruction-style data and fine-tuning LLaMA-7B (Zhang et al.,, 2023). Their models outperformed FinBERT and zero-shot ChatGPT in sentiment classification, showing improved performance in context-sensitive inputs. Wang et al., (2023) further supported this, confirming the need for instruction-tuned models in finance. Moreover, based on systematic evaluations of instruction-tuned LLMs across finance it was found that instruction tuning consistently improves performance, enabling smaller models to close the gap with larger ones (Fatemi et al.,, 2025). However, limitations remain beyond sentiment analysis, such as long-form document parsing (e.g., annual reports and 10-Ks) and valuation modelling (Li et al.,, 2023; Jadhav and Mirza,, 2025).
@@
Our work connects to research on decision-support systems, since BondBERT can be viewed as a perceptual agent module embedded within multi-agent financial architectures, where agents handle forecasting, risk assessment, and trading.
@@ 3 Experimental setup
We use two primary data sources, bond market transactional data and financial news articles. The transactional bond market data was provided by Propellant.digital, a Software as a Service (SaaS) company that has developed a platform that aggregates and standardises trading data across markets, serving as a working prototype for a consolidated tape in the bond market. [^2] https://propellant.digital/ The data was extracted from their Big Query client.api and the subset we used covers all trading days from January 2022 to January 2025. To ensure representative market movements, bonds were filtered for sufficient liquidity and volatility, and daily prices were transformed into logarithmic returns for better analysis. The final dataset includes price information for the 10 most liquid UK sovereign bonds in that period. UK corporate bonds were deliberately excluded from the analysis due to their relative illiquidity.
All bonds are issued with a notational face value (also known as “par value” or “principal”), which is the amount the issuer promises to pay to the bondholder at maturity. Throughout this work, we use the term “price” to represent trade price as a percentage of the notational value; for example, if a bond with face value £1000 trades at £1010, we record this trade as (percentage of par) price 101. Price 100 therefore indicates “trading at par”; prices above 100 indicate “trading at a premium”; and prices below 100 indicate “trading at a discount”. Prices tend not to deviate far from the par value 100, and we see this in the distribution of bond trade prices for the 10 instruments shown in Figure 1. Daily price changes also tend to be small, with most log returns close to zero, as shown in Figure 1. We have chosen 10 “liquid” bonds that are traded frequently, therefore prices react quickly to financial news. In total, the 10 bonds were traded around 30,000 times in January 2022, and trade frequency increased over the period, reaching around 100,000 monthly trades in January 2025.
@@
For reproducibility, we have open-source released all preprocessing scripts, labelling prompts, and model fine-tuning configurations on GitHub. [^4] https://github.com/Tjbarter/Sentiment-Bond-Analysis
@@ 4 Methodology
To address the absence of a bond-specific sentiment tool, we developed BondBERT by further adapting a financial domain language model for the fixed-income context. Specifically, we start from the FinBERT model introduced by Yang et al., (2020), which is domain pretrained on a large scale financial text (e.g., SEC filings and Reuters news) but not fine-tuned on labeled sentiment datasets. We fine-tuned this model on our bond-specific sentiment dataset to create BondBERT. For comparison, we use ProsusAI/FinBERT as a baseline, [^5] https://huggingface.co/ProsusAI/finbert which is widely used in prior work and is fine-tuned on the Financial PhraseBank dataset (Malo et al.,, 2014) (general financial news sentiment) but not specialised for bonds.
@@
The training corpus for BondBERT comprised approximately 17,000 international bond-related news articles spanning 2018–2021, covering macroeconomic, fiscal, and policy domains relevant to bond markets. [^6] Topics include: “macroeconomic trends”, “macroeconomic outlook”, “bond market”, “global economic growth”, “economic outlook”, “economic uncertainty”, “interest rates”, “monetary policy”, “global risk sentiment”, “financial market volatility”, “European bond market”, “energy crisis”, “global inflation”, “global GDP slowdown”, “macroeconomic report”, “bond market”, “investor sentiment”, “sanctions and trade wars”, and “geopolitical tensions”. These articles were cleaned for duplicates, irrelevant topics, and insufficient content length, then aligned to market dates. Sentiment labels were generated using GPT-4.1-nano via the OpenAI API and evaluating different prompts designed to capture the expected impact of each article on bond prices. [^7] We evaluated different GPT models, such as GPT-4.1 and GPT-4.1-mini, but due to similar results, we opted for GPT-4.1-nano, which is more computationally efficient.
@@
We used GPT-4.1-nano to efficiently generate sentiment labels, testing prompt formats until we identified one that produced consistent and financially relevant outputs. This was critical because general-purpose prompts often misinterpret fixed income market signals (e.g., interpreting “interest rate rise” as positive, even though it is negative for bonds and will lead to bond prices falling). However, we recognise that this approach introduces potential label noise and bias due to the model’s prior training data and prompt sensitivity. To mitigate this, we conducted prompt engineering experiments and cross-checked a subset of labels against known market reactions, though future work should incorporate human annotated or hybrid validation to improve model accuracy.
@@
We constructed a set of twelve representative financial news snippets, four clearly negative, four clearly positive, and four neutral, based on real bond market reporting. These examples included diverse contexts such as central bank rate hikes, inflation reports, bond issuance announcements, and fiscal policy updates. Each candidate prompt was evaluated on the basis of its alignment with expected bond price reactions and its consistency across similar examples.
@@
The best performing prompt was concise and framed the model as a bond market analyst, improving label consistency and ensuring that positive scores reflected bond price increases rather than general economic optimism. We then asked the model to assign a sentiment score between 1-1 (strongly negative for bond prices) and +1+1 (strongly positive), with 0 indicating neutral. The final prompt, which was deliberately kept concise for better accuracy, was:
@@
Using this prompt, we processed 17,000 articles for training, almost 2,000 for validation, and 2,000 for testing, in order to generate continuous sentiment scores in the range [1,1][-1,1]. These scores served as training labels for BondBERT’s supervised fine-tuning stage. We merged sentiment scores with the article text, using titles as unique keys to ensure accurate label alignment.
We initially considered further fine-tuning FinBERT on bond-specific sentiment data. However, because FinBERT’s prior training was optimised for general financial sentiment, we expected additional fine-tuning might conflict with existing scores and introduce noise rather than improve alignment with bond price movements. For this reason, we developed BondBERT as a separate model, pre-trained on financial text but fine-tuned on bond market data. Sentiment scores were categorised into three categories, negative, neutral, and positive, and stratified undersampling was applied to mitigate class imbalance by creating equal sample sizes across the categories.
@@
Table 1: BondBERT configuration parameters.
@@
Both BondBERT and FinBERT output probabilities for three sentiment classes: negative, neutral, and positive. Following a process similar to Borovkova and Zhang, (2022), we convert sentiment probabilities into a single continuous sentiment score $s\in[-1,1]$ using the Normalised Difference Index (NDI):
@@
where $s$ is the sentiment score, p(positive) is the probability that the text is positive and p(negative) is the probability that the text is negative. This transformation is particularly useful in our context because it explicitly removes the neutral class and amplifies the polarity of the sentiment signal. By focusing only on positive and negative components, NDI more directly captures information that can move the market.
@@
We used Instruct-FinGPT in its 13B parameter version, accessed through the Hugging Face inference endpoint. [^9] https://huggingface.co/FinGPT/Instruct-FinGPT-13B No LoRA adapters were attached since the model is already instruction-tuned on financial sentiment. Instruct-FinGPT expects input formatted as task-specific instructions followed by a text passage, otherwise outputs are not more reliable than zero-shot prompting. Therefore, for a fair comparison, we used the same BondBERT prompt broken into smaller sentences with guidelines to avoid chattiness:
@@
Finally, it is worth clarifying that because FinGPT and Instruct-FinGPT are not stand-alone architectures, we had to use LLaMA-2 as their underlying backbone. These models were released as adaptations of the LLaMA-2 family (7B and 13B), with LoRA adapters provided for FinGPT and instruction-tuned checkpoints for Instruct-FinGPT specifically designed for this backbone. Although newer open-source backbones such as LLaMA-3 and LLaMA-4 have been released, the FinGPT resources currently available are not compatible with them. Since our goal was to compare BondBERT with other financial models, we included FinGPT and Instruct-FinGPT in the evaluation despite their dependence on LLaMA-2.
One drawback of using GPT-based annotation is the potential risk of look-ahead bias. Since GPT models are trained on large post-event corpora, they may inadvertently leverage knowledge of subsequent market outcomes when labelling historical articles. In our current setup, we did not impose strict temporal framing in the prompt, so we cannot fully rule out this risk. However, the prompt was designed to request the immediate expected impact of each article on bond prices, which partly reduces the chance of future knowledge being incorporated. Future work should address this more rigorously by enforcing time-consistent prompting or complementing synthetic labels with human annotation. This applies not only to GPT-4.1-nano, but also to Instruct-FinGPT and any LLM-based annotator.